# Exploring Weak measurements within the Einstein-Dirac Cosmological framework

### v1

@@ III Weak measurement on the Einstein-Dirac solutions
The real part of the numerator of the weak value of $\sigma^{z}$ operator in general terms with the above conditions ($t_{\text{in}}=0$, $\Delta_{\text{in}}=0$ and $t_{\text{out}}=t_{\text{max}}$), $\sigma_{r,w}$, is given by
${w_{f}}^{3}+{w_{i}}^{3}=$
$\begin{split}&\sin(2g(t))\left(\Delta\rho_{\text{in}}^{2}\cos\left(2F\left(t,t_{\text{in}}\right)\right)+2\rho_{out_{1}}\rho_{out_{2}}\cos\left(\Delta_{\text{out}}\right)\right)+\\ &\left(\Delta\rho_{\text{out}}^{2}-2\rho_{in_{1}}\rho_{in_{2}}\right)\cos(2g(t)).\end{split}$ (128)
In terms of the preselected and postselected we have computed, it would be
$\begin{split}&\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos\left(2F\left(t,t_{\text{in}}\right)\right)\sin(2g(t))+\frac{\left(-mR_{0}^{2}+\lambda R_{\max}\right)\cos(2g(t))}{m\lambda}+\\ &\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\sin(2g(t))\sin\left(2F\left(t_{\text{out}},t\right)+\phi_{\max}\right).\end{split}$ (129)
In the same way, the imaginary part of the numerator of the weak value of $\sigma^{z}$ operator is expressed, in general term with the above conditions as,
@@
$\begin{split}&2\rho_{out_{1}}\rho_{out_{2}}\left(\Delta\rho_{\text{in}}^{2}\cos(2g(t))\sin\left(2F\left(t,t_{\text{in}}\right)+\Delta_{\text{out}}\right)+\right.\\ &\left.\rho_{in_{1}}\rho_{in_{2}}\sin(2g(t))\sin\left(\Delta_{\text{out}}\right)\right)-\Delta\rho_{\text{in}}^{2}\Delta\rho_{\text{out}}^{2}\sin(2g(t))\sin\left(2F\left(t,t_{\text{in}}\right)\right).\end{split}$ (130)
It would explicitly be given in terms of the preselected and postselected state as
$\begin{split}&-\frac{\sin\left(2F\left(t,t_{\text{in}}\right)\right)\sin(2g(t))\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}R_{\max}}{m}+\\ &\frac{1}{\lambda}\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\left(R_{0}^{2}\cos\left(2F\left(t_{\text{out}},t\right)+\phi_{\max}\right)\sin(2g(t))+\right.\\ &\left.\lambda\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos(2g(t))\cos\left(2F\left(t,t_{\text{in}}\right)-2F\left(t_{\text{out}},t\right)-\phi_{\max}\right)\right).\end{split}$ (131)
@@
$\begin{split}&2\Delta\rho_{\text{in}}^{2}\rho_{out_{1}}\rho_{out_{2}}\cos\left(2F\left(t,t_{\text{in}}\right)+\Delta_{\text{out}}\right)-2\Delta\rho_{\text{out}}^{2}\rho_{in_{1}}\rho_{in_{2}}+1.\end{split}$ (132)
It would explicitly be given by:
@@
wherein $a=2F\left(t,t_{\text{in}}\right)+\Delta_{\text{out}}$.
@@
We remind that $\Delta_{\text{out}}=(\eta_{out_{1}}-\eta_{out_{2}}+2F(t_{\text{out}},t))$. By adjusting adequately the parameters appearing in $a$, one can amplify the weak values.
@@
III.3 Berry phase from the weak measurement of a pure state.
Let us now compute the weak measurement of a pure state which is given as:
@@
For readability and simplicity we shall consider
$\theta=\varphi=\frac{\pi}{2}$, $a={\left(mt_{\text{out}}-\frac{\pi}{2}-F(t_{\text{out}},t_{\text{in}})\right)}$ and $b=2F\left(t,t_{\text{in}}\right)+\Delta_{\text{out}}$.
Then the real part of the numerator of $\hat{\Pi}_{r,w}$ will be given in general terms by
$\begin{split}&-\Delta\rho_{\text{in}}^{2}\sin\left(2F\left(t,t_{\text{in}}\right)\right)+2\rho_{out_{1}}\rho_{out_{2}}\sin\left(\Delta_{\text{out}}\right)+1.\end{split}$ (138)
In terms of the preselected and postselected state computed, it would be
$\begin{split}-\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\sin\left(2F\left(t,t_{\text{in}}\right)\right)+\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\sin\left(\Delta_{\text{out}}\right)+1\end{split}.$ (139)
The imaginary part of numerator of the $\hat{\Pi}_{r,w}$,
$1+\vec{f}.\vec{r}+\vec{r}.\vec{i}+\vec{f}.\vec{i}$, obtained is in general terms
$\begin{split}&-\Delta\rho_{\text{in}}^{2}\Delta\rho_{\text{out}}^{2}\cos\left(2F\left(t,t_{\text{in}}\right)\right)-4\rho_{in_{1}}\rho_{in_{2}}\rho_{out_{1}}\rho_{out_{2}}\cos\left(\Delta_{\text{out}}\right).\end{split}$ (140)
In terms of our particular solutions for the preselected and postselected state, it would be given by
$\begin{split}&-\frac{R_{\max}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}}{m}\cos\left(2F\left(t,t_{\text{in}}\right)\right)\\ &-\frac{R_{0}^{2}\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}}{\lambda}\cos\left(\Delta_{\text{out}}\right).\end{split}$ (141)
In the denominator, $1+\vec{f}.\vec{i}$ remains the same
$\begin{split}&2\Delta\rho_{\text{in}}^{2}\rho_{out_{1}}\rho_{out_{2}}\cos\left(b\right)-2\Delta\rho_{\text{out}}^{2}\rho_{in_{1}}\rho_{in_{2}}+1\end{split}.$ (142)
Explicitly we have
$\begin{split}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\cos\left(2F\left(t,t_{\text{in}}\right)+\Delta_{\text{out}}\right)-\frac{R_{0}^{2}R_{\max}}{\lambda m}+1\end{split}.$ (143)
@@
Let us compute now the argument of the weak value of the pure state. It is given from ([7]) as
@@
$\arctan{\left(\frac{-\frac{R_{\max}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}}{m}\cos\left(2F\left(t,t_{\text{in}}\right)\right)-\frac{R_{0}^{2}\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}}{\lambda}\cos\left(\Delta_{\text{out}}\right)}{-\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\sin\left(2F\left(t,t_{\text{in}}\right)\right)+\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\sin\left(\Delta_{\text{out}}\right)+1}\right)}.$ (145)
@@ IV CONCLUSION
Our objectives in this paper were to apply the theory of weak measurement in Einstein-Dirac system in the Friedman-Lemaître-Robertson-Walker space. We have elucidated numerous conclusions, particularly regarding the enlarged scope of weak measurements in contrast to classical measurement paradigms. Notably, our analysis revealed that the weak value of the energy-momentum tensor extends beyond what is described classical energy-momentum concepts, exhibiting unusual values as complex numbers in certain instances.
@@
Our analytical framework also demonstrates the effectiveness of weak measurements in detecting the acceleration of the Universe, leveraging the Two-State vector formalism and weak measurement techniques. Notably, we established that the scale function may be understood as an outcome of the postselection process.
We end this paper by noticing that it is possible to use this tool in cosmology. This should be the next step for some future researches.
@@ Appendix A Weak measurements of the Energy momentum
A.1 Computation of the preselected state at time t
The spinor at time $t$ is obtained by applying the evolutionary operator on a spinor defined at a specific time. The case of preselected state at time $t$ is the preselected state at time $t_{\text{in}}$ on which the unitary operator is applied. It is given by
@@ A.1 Computation of the preselected state at time t
The specific case wherein the preselected state is taken in the radiation dominated Universe with $m=0$ will yield $g(t_{\text{in}})=\frac{1}{2}\arctan(\frac{\lambda}{0*R(t_{\text{in}})})=\frac{\pi}{4}+k\pi$. For $k=0$, the matrix will be given by $U(t_{\text{in}})=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$.
@@
A.2 Computation of the postselected state at time t
The postselected state is taken from the dust dominated Universe wherein $\lambda=0$. This will yield $g(t)=\frac{1}{2}\arctan(\frac{0}{m*R(t_{\text{in}})})=0+k\pi$ This will lead, for $k=0$, to $U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$, the identity matrix.
@@ Appendix B Weak measurements of $\sigma^{z}$ operator
$\begin{split}&\alpha_{\text{out}}(t)\bar{\beta}_{\text{out}}(t)=-\left(|\alpha_{\text{out}}(t_{\text{out}})|^{2}-|\beta_{\text{out}}(t_{\text{out}})|^{2}\right)\sin(g(t))\cos(g(t))\\ &+\overline{\alpha_{\text{out}}}(t_{\text{out}})\beta_{\text{out}}(t_{\text{out}})\cos^{2}(g(t))e^{-2iF(t_{\text{out}},t)}\\ &-\overline{\beta_{\text{out}}}(t_{\text{out}})\alpha_{\text{out}}(t_{\text{out}})\sin^{2}(g(t))e^{2iF(t_{\text{out}},t)},\end{split}$ (156)
which in terms of $\alpha_{\text{out}}(t_{\text{out}})=\rho_{out_{1}}e^{i\eta_{out_{1}}}$ and $\beta_{\text{out}}(t_{\text{out}})=\rho_{out_{2}}e^{i\eta_{out_{2}}}$ will be
@@
In the same way, the preselected state vector is computed as follows $\vec{w_{i}}=$
@@
$\begin{split}&\alpha_{\text{in}}(t)\bar{\beta}_{\text{in}}(t)\\ &=\left(|\alpha_{\text{in}}(t_{\text{in}})|^{2}-|\beta_{\text{in}}(t_{\text{in}})|^{2}\right)\\ &\quad\times\Biggl{(}\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}-\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}\Biggr{)}\\ &+\alpha_{\text{in}}(t_{\text{in}})\overline{\beta}_{\text{in}}(t_{\text{in}})\\ &\quad\times\Biggl{(}\sin(2g(t))+\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}+\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}\Biggr{)}\\ &+\beta_{\text{in}}(t_{\text{in}})\overline{\alpha}_{\text{in}}(t_{\text{in}})\\ &\quad\times\Biggl{(}\sin(2g(t))-\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}-\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}\Biggr{)},\end{split}$ (159)
which in terms of $\alpha_{\text{in}}(t_{\text{in}})=\rho_{in_{1}}e^{\eta_{in_{1}}}$ and $\beta_{\text{in}}(t_{\text{in}})=\rho_{in_{2}}e^{\eta_{in_{2}}}$ will be
@@
In our case, $\vec{r}=(\begin{array}[]{ccc}0&0&1\end{array})$ to get $\sigma^{z}$. Following Kofman in [33], the vectors $\vec{f}$ and $\vec{i}$ are computed as follows
@@
In the same way, the preselected state is given by:
@@
The vector $\vec{f}$ is equal to $\frac{\vec{w_{f}}}{|\vec{w_{f}}|},\text{and }\vec{i}=\frac{\vec{w_{i}}}{|\vec{w_{i}}|}$. They represent the unit vectors on the Bloch sphere of post-selected and pre-selected state. In terms of the components, the $\sigma^{z}$ weak measurement is given by:
@@
wherein $\eta_{in_{1}}-\eta_{in_{2}}=\Delta_{in}$, $\rho_{in_{1}}^{2}-\rho_{in_{2}}^{2}=\Delta\rho^{2}_{in}$, $|\vec{w_{f}}|=\lambda^{2}-\frac{1}{4}$
The real and the imaginary part of the numerator of the weak measurement of $\sigma^{z}$ operator , $\sigma_{r,w}$, are given respectively by:
${w_{f}}^{3}+{w_{i}}^{3}=$

### v2

@@ III Weak measurement on the Einstein-Dirac solutions
The real part of the numerator of the weak value of $\sigma^{z}$ operator in terms of the pre-selected and post-selected we have computed will be given as ${w_{f}}^{3}+{w_{i}}^{3}=$
$\sin{(2g(t)}\left(\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos{(2F(t,t_{\text{in}}))}+A\cos{\left(\Delta_{\text{out}}-2F(t_{\text{out}},t)\right)}\right)$ (134)
$+\cos{(2g(t))}\left(\frac{R_{\text{max}}}{m}-\frac{R_{0}^{2}}{\lambda}\right).$
In the same way, the imaginary part of the numerator of the weak value of $\sigma^{z}$ operator in terms of the pre-selected and post-selected state is given as
@@
$\begin{split}&\sin{(2g(t))}\times\\ &\left(-\frac{R_{\text{max}}}{m}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\sin{\left(2F\left(t,t_{\text{in}}\right)\right)}+A\frac{R_{0}^{2}}{\lambda}\sin{(\Delta_{\text{out}}-2F(t_{\text{out}},t)}\right)\\ &+A\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos{(2g(t))}\sin{(\Delta_{\text{out}}+2F\left(t,t_{\text{in}}\right)+2F\left(t,t_{\text{out}}\right))}.\end{split}$ (135)
@@
wherein $a=2F\left(t,t_{\text{in}}\right)+2F\left(t,t_{\text{out}}\right)+\Delta_{\text{out}}$.
@@
We remind that $\Delta_{\text{out}}=\eta_{out_{1}}-\eta_{out_{2}}$. By adjusting adequately the parameters appearing in $a$, one can amplify the weak values.
@@
III.3 Berry phase from the weak value of a pure state.
Let us now compute the weak value of a pure state which is given as:
@@
We obtain in terms of the pre-selected and post-selected state computed for $\Phi_{in}(t_{\text{in}})=0$ and $\theta=\varphi=\frac{\pi}{2}$
$\begin{split}1+A\sin{(\Delta_{\text{out}}-2F(t_{\text{out}},t)}-\frac{R_{0}^{2}R_{\max}}{\lambda m}\\ +\left(A\cos{(a)}-\sin{(2F(t_{\text{out}},t))}\right)\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\end{split}.$ (141)
The imaginary part of numerator of the $\hat{\Pi}_{r,w}$ is given as
$\begin{split}&-\frac{R_{\max}}{m}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos\left(2F\left(t,t_{\text{in}}\right)\right)\\ &-\frac{R_{0}^{2}}{\lambda}\sqrt{N^{2}-\frac{R_{\max}^{2}}{m^{2}}}\cos{\left(2F(t_{\text{out}},t)-\Delta_{\text{out}}\right)}.\end{split}$ (142)
The denominator, $1+\vec{f}.\vec{i}$ remains the same as computed previously
@@
Let us compute now the argument of the weak value of the pure state. It is given from [8] as
@@
$\arctan{\left(\frac{-\frac{R_{\max}}{m}\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}\cos\left(2F\left(t,t_{\text{in}}\right)\right)-\frac{R_{0}^{2}}{\lambda}A\cos\left(b\right)}{1-A\sin{(b)}-\frac{R_{0}^{2}R_{\max}}{\lambda m}+\left(A\cos{(a)}-\sin{(2F(t_{\text{out}},t))}\right)\sqrt{N^{2}-\frac{R_{0}^{4}}{\lambda^{2}}}}\right)}.$ (145)
wherein $b=2F(t_{\text{out}},t)-\Delta_{\text{out}}$.
@@ IV CONCLUSION
Our objectives in this paper were to apply the theory of weak measurement in Einstein-Dirac system in the Friedman-Lemaître-Robertson-Walker space. We begun by noticing with the help of Davies [10] that all observations and measurement are weak in the quantum sense due to the nature of the processes involve. Then, after computing the two asymptotic state vector, solutions of the Einstein-Dirac equations around the Big Bang and the Universe at its maximal value of the scale factor and finding the evolutionary operator, we computed the weak values of the energy momentum tensor and two other operators.
In this work, we have elucidated numerous conclusions, particularly regarding the enlarged scope of weak measurements in contrast to classical measurement paradigms. Notably, our analysis revealed that the weak value of the energy-momentum tensor extends beyond what is described classical energy-momentum concepts, exhibiting unusual values as complex numbers in certain instances.
@@
Our analytical framework also demonstrates the effectiveness of weak measurements in detecting the acceleration of the Universe, leveraging the Two-State vector formalism and weak measurement techniques. Notably, we established that the scale function may be understood as an outcome of the post-selection process.
We end this paper by noticing that it is possible to use this tool experimentally in cosmology. This should be the next step for some future researches.
@@ Appendix A Weak measurements of the Energy momentum
A.1 Computation of the pre-selected state at time t
The spinor at time $t$ is obtained by applying the evolutionary operator on a spinor defined at a specific time. The case of pre-selected state at time $t$ is the pre-selected state at time $t_{\text{in}}$ on which the unitary operator is applied. It is given by
@@ A.1 Computation of the pre-selected state at time t
The specific case wherein the pre-selected state is taken in the radiation dominated Universe with $m=0$ will yield $g(t_{\text{in}})=\frac{1}{2}\arctan(\frac{\lambda}{0*R(t_{\text{in}})})=\frac{\pi}{4}+k\pi$. For $k=0$, the matrix will be given by $U(t_{\text{in}})=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$.
@@
A.2 Computation of the post-selected state at time t
The post-selected state is taken from the dust dominated Universe wherein $\lambda=0$. This will yield $g(t)=\frac{1}{2}\arctan(\frac{0}{m*R(t_{\text{in}})})=0+k\pi$ This will lead, for $k=0$, to $U(t_{out})=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$, the identity matrix.
@@ Appendix B Weak measurements of $\sigma^{z}$ operator
$\begin{split}&\alpha_{\text{out}}(t)\bar{\beta}_{\text{out}}(t)=-\left(|\alpha_{\text{out}}(t_{\text{out}})|^{2}-|\beta_{\text{out}}(t_{\text{out}})|^{2}\right)\sin(g(t))\cos(g(t))\\ &+\overline{\alpha_{\text{out}}}(t_{\text{out}})\beta_{\text{out}}(t_{\text{out}})\cos^{2}(g(t))e^{-2iF(t_{\text{out}},t)}\\ &-\overline{\beta_{\text{out}}}(t_{\text{out}})\alpha_{\text{out}}(t_{\text{out}})\sin^{2}(g(t))e^{2iF(t_{\text{out}},t)},\end{split}.$ (156)
In terms of $\alpha_{\text{out}}(t_{\text{out}})=\rho_{out_{1}}e^{i\eta_{out_{1}}}$ and $\beta_{\text{out}}(t_{\text{out}})=\rho_{out_{2}}e^{i\eta_{out_{2}}}$, the real part will be
@@
In the same way, the pre-selected state vector is computed as follows $\vec{w_{i}}=$
@@
$\begin{split}&\alpha_{\text{in}}(t)\bar{\beta}_{\text{in}}(t)\\ &=\left(|\alpha_{\text{in}}(t_{\text{in}})|^{2}-|\beta_{\text{in}}(t_{\text{in}})|^{2}\right)\\ &\quad\times\Biggl{(}\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}-\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}\Biggr{)}\\ &+\alpha_{\text{in}}(t_{\text{in}})\overline{\beta}_{\text{in}}(t_{\text{in}})\\ &\quad\times\Biggl{(}\sin(2g(t))+\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}+\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}\Biggr{)}\\ &+\beta_{\text{in}}(t_{\text{in}})\overline{\alpha}_{\text{in}}(t_{\text{in}})\\ &\quad\times\Biggl{(}\sin(2g(t))-\sin^{2}(g(t))e^{2iF(t,t_{\text{in}})}-\cos^{2}(g(t))e^{-2iF(t,t_{\text{in}})}\Biggr{)},\end{split}.$ (159)
In terms of $\alpha_{\text{in}}(t_{\text{in}})=\rho_{in_{1}}e^{\eta_{in_{1}}}$ and $\beta_{\text{in}}(t_{\text{in}})=\rho_{in_{2}}e^{\eta_{in_{2}}}$, the real part will be
@@
Following Kofman in [35], the vectors $\vec{f}$ and $\vec{i}$ are computed as follows
@@
In the same way, the pre-selected state is given by:
@@
wherein $\eta_{in_{1}}-\eta_{in_{2}}=\Delta_{in}$, $\rho_{in_{1}}^{2}-\rho_{in_{2}}^{2}=\Delta\rho^{2}_{in}$,
$|\vec{w_{f}}|=\lambda^{2}-\frac{1}{4}$. The vector $\vec{f}$ is equal to $\frac{\vec{w_{f}}}{|\vec{w_{f}}|},\text{and }\vec{i}=\frac{\vec{w_{i}}}{|\vec{w_{i}}|}$. They represent the unit vectors on the Bloch sphere of post-selected and pre-selected state and $\vec{r}=(\begin{array}[]{ccc}0&0&1\end{array})$. In terms of the components, the $\sigma^{z}$ weak value is given by:
The real and the imaginary part of the numerator of the weak measurement of $\sigma^{z}$ operator , $\sigma_{r,w}$, are given respectively by: ${w_{f}}^{3}+{w_{i}}^{3}=$